OpenStudy (anonymous):

suppose you buy a tv that costs $299.00 plus 6% sales tax. find the amount of sales tax andthe total bill suppose you buy a tv that costs $299.00 plus 6% sales tax. find the amount of sales tax andthe total bill @Mathematics

OpenStudy (agreene):

Tax = Cost * %Tax Tax= 299.00 * 0.06 Tax= 17.94 Total Cost = Cost + tax Total cost = 299.00+17.94 Total cost = 316.94 All in whatever currency you choose.

OpenStudy (anonymous):

17.94 316.94